OpenStudy (anonymous):

Which property justifies this statement? r(s + t) = rs + rt A. associative B. distributive C. symmetric D. transitive please help!!

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

A?

OpenStudy (skullpatrol):

Note how the r distributes as a multiplier r(s + t) = rs + rt

OpenStudy (skullpatrol):

$$\huge r(s + t) = r*s + r*t$$

OpenStudy (anonymous):

so... B?

OpenStudy (skullpatrol):

The full name of the property is "The distributive property of multiplication with respect to addition."

OpenStudy (skullpatrol):

It is a property of all real numbers.